OpenStudy (anonymous):

can any one help me please..... what are the roots of (sinx)^5+(cosx)^3=1

OpenStudy (dumbcow):

2 roots are x=0, x=pi/2

OpenStudy (anonymous):

please show your work.....thnks

OpenStudy (dumbcow):

to get the rest you have to solve an 8th degree polynomial using approximation methods

OpenStudy (anonymous):

ok..how do you start them please?

OpenStudy (dumbcow):

well think what makes sinx =0 and cosx=1 this yields x=0 well think what makes sinx =1 and cosx=0 this yields x=pi/2

OpenStudy (anonymous):

yeah correct,.... what about the polynomial?

OpenStudy (dumbcow):

ok i assigned u=sinx and v=cosx u=sinx = sqrt(1-cosx^2) = sqrt(1-v^2) from u^5 +v^3=1 u = 5th root of (1-v^3) set these equal to each other and raising to power of 10 yields (1-v^2)^5 = (1-v^3)^2 => v^2(v^8-5v^6+11v^4-10v^2-2v+5) = 0

OpenStudy (anonymous):

wow thank you so much...thats very good....

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

you are really the life saver here....lol thnk you so much again......

OpenStudy (dumbcow):

:) also depending on the directions, pi would also work

OpenStudy (dumbcow):

wait nevermind pi does not work cospi = -1

OpenStudy (anonymous):

ah ok......

OpenStudy (anonymous):

can we try also euler identities here?

OpenStudy (anonymous):

ok thnks you so much again

OpenStudy (dumbcow):

umm maybe, usually thats when you are dealing with complex numbers isn't it im not an expert on eulers

OpenStudy (anonymous):

ok for euler formula cosx=(e^ix + e^-ix)/2 sinx=(e^ix - e^-ix)/2i

OpenStudy (dumbcow):

ok if we use that we still get a messy equation (e^ix - e^-ix)^5 + 4i(e^ix +e^-ix)^3 = 32i

OpenStudy (anonymous):

thats true its too messy...lol so ill stick to polynomials lol....

OpenStudy (anonymous):

thanks again for you great great help......

OpenStudy (dumbcow):

no problem